The utopia package^{∗}

Peter Dyballa

1994/10/02

1 Introduction

This file contains the external font information needed to load the Utopia font, and a suitable package file to utilize it.

From this file all .fd files (font definition files) for the PostScript fonts with old encoding (OT1) are created, as well as those with the new encoding (T1).

2 Customization

The usual customization will be to add support for additional combinations in this file, such as a derived bold slanted font.

3 The docstrip modules

As always we begin by identifying the latest version of the files in the log file. The explicit spaces are necessary in an .fd file and the \string guards against situations where ` is active.

1 ⟨∗package⟩

 2 \ProvidesPackage{utopia}[\filedate\space\fileversion\space
 3 PSNFSS2e LaTeX2e package]
 4 \renewcommand{\rmdefault}{put}
 5 \def\bfdefault{b}

6 ⟨/package⟩

 7 % \subsection{EC-encoded Adobe Utopia-Regular}
 8 %

9 ⟨∗T1⟩

10 \typeout{File\space T1put.fd\space loading \space Adobe\space Utopia-Regular}%
11 \DeclareFontFamily{T1}{put}{}%
12 \DeclareFontShape{T1}{put}{m}{n}{<->putr8t}{}%
13 \DeclareFontShape{T1}{put}{m}{sc}{<->putrc8t}{}%
14 \DeclareFontShape{T1}{put}{m}{it}{<->putri8t}{}%

∗This file has version number v1.5i, last revised 1994/10/02.

15 \DeclareFontShape{T1}{put}{m}{sl}{<->putro8t}{}%
16 \DeclareFontShape{T1}{put}{m}{u}{<->putru8t}{}%
17 \DeclareFontShape{T1}{put}{b}{n}{<->putb8t}{}%
18 \DeclareFontShape{T1}{put}{b}{it}{<->putbi8t}{}%
19 \DeclareFontShape{T1}{put}{b}{sc}{<->putbc8t}{}%
20 \DeclareFontShape{T1}{put}{b}{sl}{<->putbo8t}{}%
21 \DeclareFontShape{T1}{put}{b}{u}{<->putbu8t}{}%
22 \DeclareFontShape{T1}{put}{bx}{n}{<->putbre8t}{}%
23 \DeclareFontShape{T1}{put}{bx}{it}{<->putbie8t}{}%
24 \DeclareFontShape{T1}{put}{bx}{sc}{<->putbce8t}{}%
25 \DeclareFontShape{T1}{put}{bx}{sl}{<->putboe8t}{}%
26 \DeclareFontShape{T1}{put}{bx}{u}{<->putbue8t}{}%
27 \DeclareFontShape{T1}{put}{mn}{n}{<->putrrn8t}{}%
28 \DeclareFontShape{T1}{put}{mn}{sc}{<->putrcn8t}{}%
29 \DeclareFontShape{T1}{put}{mn}{it}{<->putrin8t}{}%
30 \DeclareFontShape{T1}{put}{bn}{n}{<->putbrn8t}{}%
31 \DeclareFontShape{T1}{put}{bn}{it}{<->putbin8t}{}%
32 \DeclareFontShape{T1}{put}{bn}{sc}{<->putbcn8t}{}%

33 ⟨/T1⟩

34 % \subsection{Old \TeX-encoded Adobe Utopia-Regular}

35 % 36 ⟨∗OT1⟩

37 \typeout{File\space OT1put.fd\space loading \space Adobe\space Utopia-Regular}%
38 \DeclareFontFamily{OT1}{put}{}%
39 \DeclareFontShape{OT1}{put}{m}{n}{<->putr7t}{}%
40 \DeclareFontShape{OT1}{put}{m}{sc}{<->putrc7t}{}%
41 \DeclareFontShape{OT1}{put}{m}{it}{<->putri7t}{}%
42 \DeclareFontShape{OT1}{put}{m}{sl}{<->putro7t}{}%
43 \DeclareFontShape{OT1}{put}{m}{u}{<->putru7t}{}%
44 \DeclareFontShape{OT1}{put}{b}{n}{<->putb7t}{}%
45 \DeclareFontShape{OT1}{put}{b}{it}{<->putbi7t}{}%
46 \DeclareFontShape{OT1}{put}{b}{sc}{<->putbc7t}{}%
47 \DeclareFontShape{OT1}{put}{b}{sl}{<->putbo7t}{}%
48 \DeclareFontShape{OT1}{put}{b}{u}{<->putbu7t}{}%
49 \DeclareFontShape{OT1}{put}{bx}{n}{<->putbre7t}{}%
50 \DeclareFontShape{OT1}{put}{bx}{it}{<->putbie7t}{}%
51 \DeclareFontShape{OT1}{put}{bx}{sc}{<->putbce7t}{}%
52 \DeclareFontShape{OT1}{put}{bx}{sl}{<->putboe7t}{}%
53 \DeclareFontShape{OT1}{put}{bx}{u}{<->putbue7t}{}%
54 \DeclareFontShape{OT1}{put}{mn}{n}{<->putrrn7t}{}%
55 \DeclareFontShape{OT1}{put}{mn}{sc}{<->putrcn7t}{}%
56 \DeclareFontShape{OT1}{put}{mn}{it}{<->putrin7t}{}%
57 \DeclareFontShape{OT1}{put}{bn}{n}{<->putbrn7t}{}%
58 \DeclareFontShape{OT1}{put}{bn}{it}{<->putbin7t}{}%
59 \DeclareFontShape{OT1}{put}{bn}{sc}{<->putbcn7t}{}%

60 ⟨/OT1⟩